OpenStudy (anonymous):

Find the zeroes of the following cubic function: f(x)=x^3-8x-3

OpenStudy (ipwnbunnies):

Not a nice function to factor. Your best bet is to graph it and locate the zeroes, where the graph touches the x-axis, or where y-0.

OpenStudy (aum):

Have you covered rational roots theorem?

OpenStudy (aum):

|constant term| / |coefficient of x^3| = 3/1 = 3 factors of 3 are: 1 and 3 Possible roots are: -1, 1, -3, 3. Try each one and see which x value makes x^3 - 8x - 3 zero. Once you find one factor, do synthetic division. The quotient will be a quadratic. Use quadratic formula to find the other two roots.

OpenStudy (anonymous):

I'm so confused, should I try and factor it?

OpenStudy (aum):

Have you been taught rational roots theorem?

OpenStudy (anonymous):

It's been a while, but doesn't ring a bell. I think I can try graphing it and then seeing where it touches the x axis? But would it be the entire point as my answer?

OpenStudy (mayankdevnani):

Simple and with better understanding :- factor them. \[\large \bf x^3-8x-3=(x-3)(x^2+3x+1)\] So,x=3 is first answer

OpenStudy (mayankdevnani):

getting it? @Ruth100

OpenStudy (anonymous):

how do you know x = 3?

OpenStudy (aum):

If you are allowed to use a graphing calculator that will be the easiest. Note down all the x-coordinates where the graph crosses the x-axis and those are the zeros/roots of the function.

OpenStudy (anonymous):

yeah that makes sense

OpenStudy (mayankdevnani):

factorize it, \[\large \bf x^3-8x-3=0\] because we have to find zeroes or you can say roots,so we equate it to zero. then,after factoring, \[\large \bf (x-3)(x^2+3x+1)=0\] so either (x-3)=0 OR \(\large \bf x^2+3x+1\)=0 so our first answer is x-3=0.So, x=3

OpenStudy (mayankdevnani):

understood? and if you have any doubt,put x=3 in your equation,if it turn into 0,then x=3 is your 1st zero or root

OpenStudy (mayankdevnani):

@Ruth100

OpenStudy (aum):

@mayankdevnani The question that Ruth had before was HOW did you factor out the (x-3). What were the steps?

OpenStudy (mayankdevnani):

oh!!..

OpenStudy (mayankdevnani):

its so simple, @ruth100 , first i do trial method, first put x=1,if they satisfy the equation to 0, then 1 is a root.but its not then,i put x=2,it also doesn't satisfy and finally,i put x=3,it satisfies.So x=3 is our first root. then next,i divide the equation by x-3 to get another equation or you can say by factorizing it,i got a new equation and then simply it and get other two new roots. If we divide, \[\large \bf \frac{x^3-8x-3}{x-3}=x^2+3x+1\]

OpenStudy (mayankdevnani):

that is my method to factorize.

OpenStudy (anonymous):

I got y= -3 X= 3 And I can't get the third point. I graphed it.

OpenStudy (mayankdevnani):

you can't..its real but you can't graph it. to find other root 2 roots :- \[\large \bf x^2+x+1=0\] use quadratic formula

OpenStudy (mayankdevnani):

correction :-

OpenStudy (anonymous):

I can still use the quadratic formula to find the other roots?

OpenStudy (mayankdevnani):

yeah

OpenStudy (mayankdevnani):

use quadratic formula to find other 2 roots

OpenStudy (anonymous):

Would I use the original given function?

OpenStudy (anonymous):

I got 8.35 And -0.35

OpenStudy (anonymous):

It seems wrong?

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

So I did my calculations wrong? I'm really confused